Hydrogen cyanide, HCN, is a weak acid with a dissociation constant of 4.8 × 10–10.

Calculate the percent dissociation, the pH, and the OH– concentration of a 0.15 M

solution of HCN in water.


1
Expert's answer
2021-03-04T06:34:11-0500

HCN "\leftrightarrow" H+ + CN-

0.15____--___--

0.15-x____x___x

"\frac{x^2}{0.15-x}" = 4.8×10-10


let suppose 0.15-x "\sim" 0.15

"\frac{x^2}{0.15}" = 4.8×10-10


x2 = 4.8×10-10 ×0.15

x2 = 0.72×10-10 = 72×10-12

x = 8.48 × 10-6 M


pH = -logoX

pH = 5.09


%dissociation =( 8.48×10-6/0.15 )×100

= 56.5×10-4 %